# for Medical Imaging Foundation Models

### v2

@@ 1 Introduction
Foundation Model (FM) facilitated medical image analysis is playing a pivotal role in healthcare [2, 3]. These models, which leverage large-scale pretraining and fine-tuning [6], have demonstrated remarkable capabilities in various medical imaging tasks, including classification [69, 41] and segmentation [55, 39]. As the use of FMs proliferates in medical imaging, addressing the challenges of evaluating and ensuring their fairness and utility becomes increasingly critical [28, 24, 77], where biases in model performance can result in significant disparities in patient care and outcomes.
@@ 3.1 Datasets
The following 17 publicly avaliable datasets are included in FairMedFM to evaluate fairness in FMs in medical imaging: CheXpert [22], MIMIC-CXR [25], HAM10000 (CLS) [61], FairVLMed10k [36], GF3300 [37], PAPILA [31], BRSET [42], HAM10000 (SEG) [61], TUSC [53], FairSeg [60], Montgomery County X-ray [7], KiTS 2023 [18], CANDI [27], IRCADb [58], SPIDER [62]. These datasets vary in the following aspects: (1) Task type: classification and segmentation; (2) Dimension: 2D and 3D; (3) Modality: OCT, X-ray, CT, MRI, Ultrasound, Fundus, OCT, and dermatology; (4) Body part: brain, eyes, skin, thyroid, chest, liver, kidney, and spine; (5) Number of classes: ranging from 2 to 15; (6) Number of samples: ranging from 20 to more than 350k; (7) Sensitive attribute: sex, age, race, preferred language, skin tone, etc. (8) SA skewness (Male : Female): ranging from 0.19 to 1.67. The details of these datasets can be found in the Appendix.
@@ 3.4 Evaluation Metrics Taxnonmy
Group fairness is evaluated from four aspects following common practice [8]: (1) Outcome-consistency fairness, which evaluates discrepancies in the model’s performance (e.g., accuracy, components of confusion matrix, etc.) between different sensitive groups. In classification, we include delta AUC ($\text{AUC}_{\Delta}$), which is measured as the maximum AUC gap among subgroups; and Equalized Odds (EqOdds), which measures the differences in true positive and false positive rates between advantaged and disadvantaged groups. In segmentation, we include delta DSC ($\text{DSC}_{\Delta}$), which assesses if both groups receive approximately equal predictive performance; and DSC skewness ($\text{DSC}_{\text{Skew}}$), which measures the degree of skewness of DSC between advantaged and disadvantaged groups. (2) Predictive alignment fairness, which focuses on the alignment between predicted probabilities and actual outcomes. It ensures that predicted scores accurately reflect true likelihoods, providing a reliable basis for decision-making across different groups. We report the expected calibration error gap ($\text{ECE}_{\Delta}$) in classification, where a high value indicates an optimal decision threshold; (3) Positive-parity fairness, which ensures that the positive classification rate is equal for both unprivileged and privileged groups, preventing any group from being overlooked. We note that this is an optional evaluative aspect that may not be applicable to all scenarios. For example, positive parity is compromised in diagnosing glaucoma, where morbidity rates differ between males and females; (4) Representation fairness, which evaluates fairness from the aspects of feature representation learned in the latent space, by estimating either group-wise feature separability among subgroups. We report the last two metrics in the Appendix.
@@
The mathematical definitions of the above metrics are presented in the Appendix.
@@ 4 Results
In this section, we highlight the representative observations and takeaways from benchmarking the fairness of FMs on image classification and segmentation tasks utilizing FairMedFM framework. We choose to present the fairness results concerning sex since it is the most common SA shared across datasets. However, our method supports a broader range of SAs as listed in the Tab. 1. We direct readers to the Appendix for additional results on more SAs and extensive evaluations, which corroborate the observations discussed in the main text. We also present the evaluation for positive-parity-fainres, representation-fairness, and statistics test in Appendix.
@@
Careful selection and use of FMs are needed for ensuring a good fairness-utility trade-off. These pervasive biases challenge the fairness-utility tradeoff of FMs in medical applications. We observe that the fairness-utility trade-off in these tasks is influenced not only by the choice of FMs but also by how they are used as shown in Fig. 4. We report trade-off scores $\text{AUC}_{\text{ES}}$ and $\text{DSC}_{\text{ES}}$, for each datasets on the selected models for both tasks respectively. Compared with CLIP-ZS models, a simple adaptation, CLIP-Adapt, has proven effective in significantly boosting the fairness-utility trade-off in medical applications. Further, we evaluate the effects in segmentation tasks on the choice of prompts (including center, rand, rands, and bbox), and the types of SegFMs (including 2D General-SegFMs and 2D Med-SegFMs). Compared to center and rand, models using rands and bbox tend to be fairer across different datasets. This might be due to the tighter constraints on the segmentation provided by rands and bbox. Regarding models, General-SegFMs achieve better fairness-utility trade-off than Med-SegFMs.
@@
Consistent disparities in SA occur across various FMs on the same dataset. The performance of FMs shows dataset-specific biases, favoring one category of the given SA over the other, depending on the dataset. We present the four datasets in classification tasks and use sex as an example, presented in Fig. 5. The BREST and MIMIC-CXR dataset shows a higher performance for female data compared to male data across various models. Conversely, the FairVLMed10k and GF3300 dataset indicates better performance for male data. More results on other SAs and segmentation tasks can be found in the Appendix.
@@ C.3 Fairness-utility Tradeoff
The fairness-utility tradeoff pursues a balance between fairness and utility. Following [37, 60], we use Equity Scaling measurements of AUC ($\text{AUC}_{\text{ES}}$) and DSC ($\text{DSC}_{\text{ES}}$) for classification and segmentation, respectively. We give the formula for AUC as an example and that of DSC is similar.
@@
where $\overline{\text{AUC}}$ is the average AUC over all data samples, and AUCΔ is the standard deviation of AUC across all subgroups defined by sensitive attributes.
@@ E.2 More Observations
Representation fairness. The t-SNE of 2D SegFMs on the four 2D datasets are shown in Fig. LABEL:. Here we use sex as an example. As shown in Fig. 16, compared to the TUSC dataset and FairSeg dataset, the latent space of the HAM10000 dataset is more separatable. This is roughly aligned with the results in Fig. 3, where the $\text{DSC}_{\Delta}$ of the HAM10000 dataset are larger than the rest datasets. This finding provides potential for us to mitigate unfairness for SegFMs by manipulating the latent space, which has been explored in APPLE [67]. On the other hand, considering that the group-wise separablity is similar acrosss different SegFMs, these four tasks might suffer more unfairness from the data than from the model.

### v3

@@ 1 Introduction
Foundation Model (FM) facilitated medical image analysis is playing a pivotal role in healthcare [2, 3, 26]. These models, which leverage large-scale pretraining and fine-tuning [6], have demonstrated remarkable capabilities in various medical imaging tasks, including classification [71, 43] and segmentation [57, 41]. As the use of FMs proliferates in medical imaging, addressing the challenges of evaluating and ensuring their fairness and utility becomes increasingly critical [30, 24, 79], where biases in model performance can result in significant disparities in patient care and outcomes.
@@ 3.1 Datasets
The following 17 publicly avaliable datasets are included in FairMedFM to evaluate fairness in FMs in medical imaging: CheXpert [22], MIMIC-CXR [27], HAM10000 (CLS) [63], FairVLMed10k [38], GF3300 [39], PAPILA [33], BRSET [44], HAM10000 (SEG) [63], TUSC [55], FairSeg [62], Montgomery County X-ray [7], KiTS 2023 [18], CANDI [29], IRCADb [60], SPIDER [64]. These datasets vary in the following aspects: (1) Task type: classification and segmentation; (2) Dimension: 2D and 3D; (3) Modality: OCT, X-ray, CT, MRI, Ultrasound, Fundus, OCT, and dermatology; (4) Body part: brain, eyes, skin, thyroid, chest, liver, kidney, and spine; (5) Number of classes: ranging from 2 to 15; (6) Number of samples: ranging from 20 to more than 350k; (7) Sensitive attribute: sex, age, race, preferred language, skin tone, etc. (8) SA skewness (Male : Female): ranging from 0.19 to 1.67. The details of these datasets can be found in the Appendix B.
@@ 3.4 Evaluation Metrics Taxnonmy
Group fairness is evaluated from four aspects following common practice [8]: (1) Outcome-consistency fairness, which evaluates discrepancies in the model’s performance (e.g., accuracy, components of confusion matrix, etc.) between different sensitive groups. In classification, we include delta AUC ($\text{AUC}_{\Delta}$), which is measured as the maximum AUC gap among subgroups; and Equalized Odds (EqOdds), which measures the differences in true positive and false positive rates between advantaged and disadvantaged groups. In segmentation, we include delta DSC ($\text{DSC}_{\Delta}$), which assesses if both groups receive approximately equal predictive performance; and DSC skewness ($\text{DSC}_{\text{Skew}}$), which measures the degree of skewness of DSC between advantaged and disadvantaged groups. (2) Predictive alignment fairness, which focuses on the alignment between predicted probabilities and actual outcomes. It ensures that predicted scores accurately reflect true likelihoods, providing a reliable basis for decision-making across different groups. We report the expected calibration error gap ($\text{ECE}_{\Delta}$) in classification, where a high value indicates an optimal decision threshold; (3) Positive-parity fairness, which ensures that the positive classification rate is equal for both unprivileged and privileged groups, preventing any group from being overlooked. We note that this is an optional evaluative aspect that may not be applicable to all scenarios. For example, positive parity is compromised in diagnosing glaucoma, where morbidity rates differ between males and females; (4) Representation fairness, which evaluates fairness from the aspects of feature representation learned in the latent space, by estimating either group-wise feature separability among subgroups. We report the last two metrics in the Appendix E.2.
@@
The mathematical definitions of the above metrics are presented in the Appendix C.
@@ 4 Results
In this section, we highlight the representative observations and takeaways from benchmarking the fairness of FMs on image classification and segmentation tasks utilizing FairMedFM framework. We choose to present the fairness results concerning sex since it is the most common SA shared across datasets. However, our method supports a broader range of SAs as listed in the Tab. 1. We direct readers to Appendix E.1 for additional results on more SAs and extensive evaluations, which corroborate the observations discussed in the main text. We also present the evaluation for positive-parity-fairness, representation-fairness, and statistics test in Appendix E.2.
@@
Careful selection and use of FMs are needed for ensuring a good fairness-utility trade-off. These pervasive biases challenge the fairness-utility tradeoff of FMs in medical applications. We observe that the fairness-utility trade-off in these tasks is influenced not only by the choice of FMs but also by how they are used as shown in Fig. 4. We report trade-off scores $\text{AUC}_{\text{ES}}$ and $\text{DSC}_{\text{ES}}$, for each datasets on the selected models for both tasks respectively. Compared with CLIP-ZS models, a simple adaptation, CLIP-Adapt, has proven effective in significantly boosting the fairness-utility trade-off in medical applications. Further, we evaluate the effects of segmentation tasks on the choice of prompts (including center, rand, rands, and bbox), and the types of SegFMs (including 2D General-SegFMs and 2D Med-SegFMs). Compared to center and rand, models using rands and bbox tend to be fairer across different datasets. This might be due to the tighter constraints on the segmentation provided by rands and bbox. Regarding models, General-SegFMs achieve a better fairness-utility trade-off than Med-SegFMs.
Consistent disparities in SA occur across various FMs on the same dataset. The performance of FMs shows dataset-specific biases, favoring one category of the given SA over the other, depending on the dataset. We present the four datasets in classification tasks and use sex as an example, presented in Fig. 5. The BREST and MIMIC-CXR dataset shows a higher performance for females compared to males across various models. Conversely, the FairVLMed10k and GF3300 dataset indicates better performance for males. Results on other SAs and segmentation tasks can be found in Appendix E.1.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
Y. Zhong and Q. Dou are supported by the Research Grants Council of Hong Kong Special Administrative Region, China (Project No. T45-401/22-N). Z. Xu, O. Yao, and S.K. Zhouare supported in part by the Natural Science Foundation of China under Grant 62271465, SuzhouBasic Research Program under Grant SYG202338, and Open Fund Project of Guangdong Academy of Medical Sciences, China (No. YKY-KF202206). R. Jin and X. Li are supported by the UBC Advanced Research Computing and Digital Research Alliance of Canada. We thank the reviewers for their insightful feedback and comments.
@@ C.3 Fairness-utility Tradeoff
The fairness-utility tradeoff pursues a balance between fairness and utility. Following [39, 62], we use Equity Scaling measurements of AUC ($\text{AUC}_{\text{ES}}$) and DSC ($\text{DSC}_{\text{ES}}$) for classification and segmentation, respectively. The equations are as follows.
@@
$\text{DSC}_{\text{ES}}$ $=\frac{\overline{\text{DSC}}}{1+\text{DSC}_{\Delta}}$ (3)
where $\overline{\text{AUC}}$ and $\overline{\text{DSC}}$ are the average AUC and DSC over all data samples. AUCΔ and DSCΔ are the standard deviation of AUC and DSC across all subgroups defined by sensitive attributes, respectively.
@@ E.2 More Observations
Representation fairness. The t-SNE of 2D SegFMs on the four 2D datasets are shown in Fig. 16. Here we use sex as an example. As shown in Fig. 16, compared to the TUSC dataset and FairSeg dataset, the latent space of the HAM10000 dataset is more separatable. This is roughly aligned with the results in Fig. 3, where the $\text{DSC}_{\Delta}$ of the HAM10000 dataset are larger than the rest datasets. This finding provides potential for us to mitigate unfairness for SegFMs by manipulating the latent space, which has been explored in APPLE [69]. On the other hand, considering that the group-wise separablity is similar acrosss different SegFMs, these four tasks might suffer more unfairness from the data than from the model.